2 to the power of 256 when divided by 17 gives the rem.----??? Pls explain!!!t

Grade:9

1 Answers

Chetan Mandayam Nayakar
12 years ago

2256=24*64=1664=(17-1)64=∑(r=0 to r=63) 64Cr1764-r(-1)r+1=17m+1, where m is a positive integer. Therefore, remainder=1